OpenStudy (anonymous):

Algebra 2 please help! Part 1: Is the following equation true? (2 points) Part 2: Use complete sentences to explain the properties used in making your decision. (6 points) 4 log 3 x + one–half log 3 y – 2 log 3 z = log 3 x to the fourth power times square root of y all over z–squared.

OpenStudy (imstuck):

The rules are as follows: \[\log _{a}(x \times y)=\log _{a}(x) + \log _{a}(y)\]AND\[\log _{a}(x)-\log _{a}(y)=\log _{a}\frac{ x }{ y }\]So let's see what we can do with that.

OpenStudy (anonymous):

I really don't understand how to apply this I've been looking at the rules but I don't udneratand

zepdrix (zepdrix):

\[\Large\rm 4\log_3 x + \frac{1}{2}\log_3 y – 2\log_3 z = \log_3 \left(\frac{x^4\sqrt y}{z^2}\right)\]This is what we're trying to determine, yes? Did I type that correctly?

OpenStudy (imstuck):

Ok, there's another one that deals with exponents, too. Then we will work on solving it.\[\log _{a}(x)^{p}=p \times \log _{a}(x)\]So let's go through this one thing at a time, if we can.

OpenStudy (anonymous):

yes @zepdrix

OpenStudy (imstuck):

The 4 is out front because, according to our rules, it was x's exponent. So the first part is \[\log _{3}(x)^{4}\]

OpenStudy (imstuck):

The 1/2 in front of the second term was at one time\[\log _{3}(y)^{\frac{ 1 }{ 2 }}\]which is also \[\log _{3}(\sqrt{y})\]

OpenStudy (imstuck):

So far we have\[\log _{3}x ^{4}+\log _{3}\sqrt{y}\]The addition sign in between them tells us that to simplify they are to be multiplied so\[\log _{3}(x \times \sqrt{y})\]

OpenStudy (imstuck):

Oops that's

OpenStudy (imstuck):

\[\log _{3}(x ^{4}\times \sqrt{y})\]

OpenStudy (imstuck):

Now to deal with the 2 log base 3 of z...the 2 was once an exponent, according to our rules. So put it back there such that\[\log _{3}z ^{2}\]

OpenStudy (imstuck):

Now we are ready to finish it off. The subtraction sign in between tells us that this is the denominator of a fraction. So the whole thing is, according to our rules of logs with the same base, is\[\frac{ \log _{3}(x ^{4}\times \sqrt{y}) }{ \log _{3}z ^{2} }\]

OpenStudy (anonymous):

what about the other two terms?

OpenStudy (anonymous):

do you do nothing with them?

OpenStudy (anonymous):

I mean the last one sorry

OpenStudy (anonymous):

@IMStuck @zepdrix I really need hep finishing this please help

zepdrix (zepdrix):

whut? :U

OpenStudy (anonymous):

can you please help finish off where he left ?

zepdrix (zepdrix):

I don't see where you left off.. around here?\[\Large\rm \log_3\left(x^4 \sqrt{y}\right) – 2\log_3 z = \log_3 \left(\frac{x^4\sqrt y}{z^2}\right)\]

zepdrix (zepdrix):

\[\Large\rm \log_3\left(x^4 \sqrt{y}\right) – \log_3 z^2 = \log_3 \left(\frac{x^4\sqrt y}{z^2}\right)\]

OpenStudy (anonymous):

no he last equation he put in log v3 (x^4 x sqrt y)/log v3 z^2

zepdrix (zepdrix):

\[\Large\rm \log_3\left(\frac{x^4 \sqrt{y}}{z^2}\right)= \log_3 \left(\frac{x^4\sqrt y}{z^2}\right)\]Here....? Then the problem is finished. See how you've shown that both sides are the same?

OpenStudy (anonymous):

how did you get that?

zepdrix (zepdrix):

I don't know why he wrote log_3 z^2 in the bottom, that must've been a typo.

zepdrix (zepdrix):

This is our log rule:\[\Large\rm \color{royalblue}{\log(a)-\log(b)=\log\left(\frac{a}{b}\right)}\]Not this:\[\Large\rm \color{red}{\log(a)-\log(b)\ne\frac{\log(a)}{\log(b)}}\]

zepdrix (zepdrix):

Maybe color will help with the rule, hold on one sec :d

zepdrix (zepdrix):

\[\Large\rm \log_3\left(\color{orangered}{a}\right)-\log_3\left(\color{royalblue}{b}\right)=\log_3\left(\frac{\color{orangered}{a}}{\color{royalblue}{b}}\right)\]Applying this to our problem:\[\Large\rm \log_3\left(\color{orangered}{x^4 \sqrt y}\right)-\log_3\left(\color{royalblue}{z^2}\right)=\log_3\left(\frac{\color{orangered}{x^4 \sqrt y}}{\color{royalblue}{z^2}}\right)\]Understand how the rule gets applied?

OpenStudy (anonymous):

why is a x4 sqrt 7 and why is b z2

zepdrix (zepdrix):

Because the subtraction is in front of the b term.

zepdrix (zepdrix):

The one that has the subtraction is the one that ends up in the denominator.

OpenStudy (anonymous):

okay i kind of get it lol thank you!

zepdrix (zepdrix):

c: math is fun! don't give up!! lol

OpenStudy (anonymous):

a bit late lol

zepdrix (zepdrix):

haha XD

OpenStudy (imstuck):

I m so sorry......my computer totally crashed! Im glad you got the help you needed to finish, though. Did not mean to leave you hanging there!